# Anomaly-induced vanishing of brane partition functions

### v1

@@ 2.2 Anomaly and the mapping cylinder
where $\mathcal{A}[\check{A},\check{\lambda}]$ is known as the ’t Hooft anomaly tHooft:1979rat . For simplicity, we have assumed here that the partition function is a section of a line bundle $\mathcal{L}$ (i.e. a vector bundle with rank $1$) over the moduli space of gauge fields, $\mathfrak{A}\simeq\check{Z}^{p+2}(X^{d})$. Equivalently, $\mathcal{A}[\check{A},\check{\lambda}]$ can be realised in terms of an invertible field theory on a $(d+1)$-dimensional bulk $Y^{d+1}$ such that $\partial Y^{d+1}=X^{d}$. More generally, we can consider replacing the line bundle with a vector bundle of higher rank, but we will not study such cases in our work.
When $\mathcal{A}[\check{A},\check{\lambda}]\notin\mathbb{Z}$, one cannot uniquely define $\mathcal{Z}[\check{A}]$ for a given differential cohomology class $\check{A}\in\check{H}^{p+2}(X^{d})$. This means that the partition function is not well-defined over the space of gauge fields modulo gauge transformations, $\mathfrak{A}/\mathfrak{G}\simeq\check{H}^{p+2}(X^{d})$. Hence, the anomaly can be interpreted as an obstruction to gauging the global symmetry, i.e. summing over all inequivalent classes of background gauge fields $\check{A}\in\check{H}^{p+2}(X^{d})$ in the partition function. (If we insist in doing such a sum, the result will vanish, but this is a different kind of vanishing to the one in this paper, which does not involve gauging of the symmetry and happens only for specific background field configurations.) We stress that this notion of an anomaly applies not only to a $p$-form $\mathrm{U}(1)$ symmetry, but also to when the symmetry is any group, e.g. non-Abelian or discrete Hopkins:2002rd ; Hsieh:2020jpj , and even when it is a higher group, as long as we define the suitable differential cocycles $\check{A}$ accordingly. In the latter case, $\mathcal{A}[\check{A},\check{\lambda}]$ is usually known as a mixed anomaly between the different levels of the associated Postnikov tower.
Geometrically, the relation (26) can be reinterpreted as a parallel transport of $\mathcal{Z}[\check{A}]$ over the mapping cylinder $X^{d}\times I$, where $I=[0,1]$ is the unit interval. More explicitly, we formally extend the gauge field to $X^{d}\times I$ such that it interpolates between two different gauge fields $\check{A},\check{A}^{\prime}\in\check{Z}^{p+2}(X^{d})$ on the two ends (we will soon impose that $A$ and $A^{\prime}$ are related by a gauge transformation), i.e.
@@
where $\check{\mathbb{A}}\in\check{C}^{p+2}(X^{d}\times I)$ is a “differential cochain” represented by the triplet, [^6] The combination $(1-t)\mathsf{C}_{p+2}+t\mathsf{C}_{p+2}^{\prime}$ is generally not an integral cochain in $C^{p+2}(X^{d}\times I;\mathbb{Z})$ for arbitrary $t\in[0,1]$, so $\check{\mathbb{A}}$, and hence the bulk gauge field $\underline{\check{A}}$, are not strictly differential cochains as we defined in (11). However, the rest of our argument does not depend on this subtlety.
@@
for some $\check{a}=(\mathsf{c}_{p+1},a_{p},f_{p+1})\in\check{C}^{p+1}(X^{d})$. [^7] To be precise, we should send both $\check{a}$ and $\check{t}$ to differential cochains in $\check{C}^{\ast}(X^{d}\times I)$ via the inclusion maps $\iota_{X}:X^{d}\hookrightarrow X^{d}\times I$ and $\iota_{I}:I\hookrightarrow X^{d}\times I$, in order to make sense of their product. When one sets $\check{a}=0$, the differential cochain $\underline{\check{A}}=\check{\mathbb{A}}\in\check{C}^{p+2}(X^{d})$ becomes “pure-shift” 2008arXiv0810.4935S .
@@
In any given QFT, we shall define the anomaly $\mathcal{A}[\check{A},\check{\lambda},\check{a}]\in\mathbb{R}$ in (26) as the evaluation of some rational polynomial $\mathcal{P}\in\mathbb{Q}[\check{A},\check{\lambda},\check{a},\check{t}]$ over the mapping cylinder $X^{d}\times I$. Recall that the bulk gauge field $\underline{\check{A}}\in\check{Z}^{p+2}(X^{d}\times I)$ is defined solely by the constraints,
@@ 2.3 Vanishing theorem for the partition function
for some choice of $\check{\lambda}\in\check{Z}_{\text{flat}}^{p+1}(X^{d})$ now regarded instead as an auxiliary gauge field. Since $\check{a}\in\check{Z}_{\text{flat}}^{p+1}(X^{d})$ is closed under the differential $d$ by construction, we have $\check{A}-d\check{a}=\check{A}$. In other words, the space of flat differential cocycles $\check{Z}_{\text{flat}}^{p+1}(X^{d})$ can be interpreted as the space of rigid gauge transformations that does not change the gauge field $\check{A}\in\check{Z}^{p+1}(X^{d})$. [^9] In the terminology of Diaconescu:2003bm ; Moore:2004jv ; Belov:2004ht ; Belov:2006jd ; Freed:2006yc , these were referred to as global gauge transformations. More explicitly, under such a transformation, we find
@@
As far as $\mathcal{A}[\check{A},0,\check{a}]$ is concerned, it will be convenient for us to carry out computations on the mapping torus $X^{d}\times S^{1}$ instead by gluing the two ends of the interval $I$, on which the bulk gauge field $\iota_{X\times\{0,1\}}^{\ast}\,\underline{\check{A}}=\check{A}\in\check{Z}^{p+2}(X^{d})$ becomes identical when pulled back to $X^{d}$. In fact, $\mathcal{A}[\check{A},0,\check{a}]$ can now be interpreted as a holonomy of the polynomial $\mathcal{P}$ over $X^{d}\times S^{1}$. The practical advantage of doing so is that the mapping torus has no boundary, so if needed $\mathcal{P}$ can be further promoted to an invariant polynomial having support over a bounding manifold $Z^{d+2}$ such that $\partial Z^{d+2}=X^{d}\times S^{1}$. For example, we can take $Z^{d+2}=X^{d}\times D^{2}$, where $D^{2}$ is the solid 2-disk with $\partial D^{2}=S^{1}$. More precisely, we would like to construct a differential cocycle $\check{\mathcal{P}}\in\check{H}^{d+2}(Z^{d+2})$ whose holonomy is given by $\chi(Z^{d+2})=\exp(2\pi i\int_{X^{d}\times S^{1}}\mathcal{P})=\mathcal{A}[\check{A},0,\check{a}]$. [^10] In general, such an extension exists if and only if the pair $(X^{d}\times S^{1},\mathsf{C}_{p+2})$ is trivial as an element of the bordism group $\Omega_{d+1}(K(\mathbb{Z},p+2))$, where $K(\mathbb{Z},p+2)$ is an Eilenberg-MacLane space whose only non-trivial homotopy group is $\pi_{p+2}=\mathbb{Z}$. The bordism group should be replaced by its suitable variants if we wish the extension to preserve some extra tangential structures. For instance, it is common to require the extension to preserve the Spin structure, in which case $S^{1}$ is null-bordant if and only if it has a Neveu-Schwarz, i.e. bounding, Spin structure. See Figure 1 for an illustration.
@@
For non-flat $\check{a}\in\check{Z}^{p+1}(X^{d})$, there is no obvious interpretation in terms of gauge transformations of the background gauge field $\check{A}\in\check{Z}^{p+2}(X^{d})$ in the $d$-dimensional QFT on $X^{d}$. One may naïvely try to use the correspondence (41), and regard $\check{a}$ as a local gauge transformation of $\check{A}$. However, this is not consistent with the fact that, by definition, a gauge equivalence is one that leaves the holonomy $\chi$ of a differential cocycle (in this case the anomaly polynomial $\mathcal{P}$) invariant, which corresponds to differential cochains, of one degree lower, that have vanishing curvature. This contradicts the assumption that $\check{a}$ is not flat. Henceforth, we restrict the auxiliary gauge field $\check{a}\in\check{Z}^{p+1}(X^{d})$ to be flat. [^11] See Diaconescu:2003bm ; Belov:2004ht for a related discussion of the more general case when $\check{a}$ is not flat. As a side remark, the requirement of $\check{a}$ being flat formally amounts to imposing a weight filtration of a differential function, here taken to be the bulk gauge field $\underline{\check{A}}$, as defined by Hopkins:2002rd (see also Diaconescu:2003bm for a concise review). However, we will not make use of this language for the purpose of our work.
@@
In contrast, the latter arises when one simply sits at some fixed basepoint $\check{A}\in\mathfrak{A}$, so a non-trivial phase $e^{2\pi i\mathcal{A}[\check{A},0,\check{a}]}$ signifies that the section $\mathcal{Z}[\check{A}]$ itself has to vanish at this point, as depicted in Figure 2. This happens because $\mathfrak{G}$ does not act freely, and its fixed points correspond precisely to the space of rigid gauge transformations, $\check{Z}_{\text{flat}}^{p+1}(X^{d})$. [^12] It is perhaps better to rephrase this picture in the language of moduli stacks. Loosely speaking, we should replace the moduli space with a lasagna-like object where each layer is a copy of $\mathfrak{A}$, then a rigid gauge transformation should correspond to a non-trivial path across layers while fixing the basepoint $\check{A}$ on each copy. This implies that the partition function $\mathcal{Z}[\check{A}]$, despite the notation, depends not only on the background gauge field, but also secretly on its automorphisms. A more detailed investigation on this perspective is left to future work. We also thank Victor Carmona for a useful discussion on this issue. For this reason, we will hereafter call $\mathcal{A}[\check{A},0,\check{a}]$ the basepoint anomaly. See Seiberg:2018ntt ; Meynet:2025zem for a similar discussion in the context of anomalies of duality groups, in which case the “gauge field” is the coupling constant $\tau$, and $\mathfrak{A}/\mathfrak{G}$ becomes $\mathbb{H}/\text{SL}(2,\mathbb{Z})$, where $\mathbb{H}$ denotes the upper-half plane of $\mathbb{C}$.
2.4 Relation to quantum Gauss law and SymTFT
For general $\mathcal{A}[\check{A},\check{\lambda},\check{a}]$, compatibility with composition requires
@@ 2.4 Relation to quantum Gauss law and SymTFT
If the basepoint anomaly $\mathcal{A}[\check{A},0,\check{a}]$ is a non-trivial element in $\mathbb{R}/\mathbb{Z}$, then $\mathcal{A}[\check{A},\check{\lambda},\check{a}]$ is also non-vanishing in general. In other words, for the relation $\mathcal{Z}[\check{A}-d\check{\lambda}]=e^{2\pi i\mathcal{A}[\check{A},\check{\lambda},\check{a}]}\mathcal{Z}[\check{A}]$ to be consistent for arbitrary rigid gauge transformations $\check{a}$, the vanishing of $\mathcal{Z}[\check{A}]$ implies the same for $\mathcal{Z}[\check{A}-d\check{\lambda}]$.
In the context where the symmetry in question is not a global symmetry but a gauge symmetry of the QFT, our analysis reproduces a previous result in the literature concerning the gauge invariance of physical wavefunctions Diaconescu:2003bm ; Moore:2004jv ; Belov:2004ht ; Belov:2006jd ; Freed:2006yc (see also Hosseini:2025oka for a recent study of anomaly cancellation in Type I string theory). Here, assuming $\mathcal{L}$ is lifted to a line bundle over $\mathfrak{A}/\mathfrak{G}$, we have $\mathcal{Z}[\check{A}-d\check{\lambda}]=\mathcal{Z}[\check{A}]$ by definition. Suppose we first neglect the auxiliary gauge field $\check{a}\in\check{Z}_{\text{flat}}^{p+1}(X^{d})$, then we demand
@@
which is known as the classical Gauss law. [^14] At a semi-classical level, Gauss law constraints can be used to construct topological operators generating the respective global symmetries. See e.g. Apruzzi:2022rei ; Choi:2022fgx ; Bah:2023ymy for some recent applications to non-invertible symmetries. On the contrary, one should also impose that the wavefunction is invariant under the action of non-trivial automorphisms $\check{a}$ of $\check{A}$. This further refines the constraint above, i.e.
@@
More precisely, despite giving analogous conclusions, there is a subtle conceptual difference between Diaconescu:2003bm ; Moore:2004jv ; Belov:2004ht ; Belov:2006jd ; Freed:2006yc and our approach. In the former, one is interested in wavefunctions $\Psi(\check{A}|_{W})$ obtained after performing a Hamiltonian quantisation on the theory. This effectively treats $X^{d}$ as a manifold with boundary $\partial X^{d}=W^{d-1}$, such that the tangent bundle near the “time-slice” $W^{d-1}$ decomposes as $TX^{d}|_{W}\cong TW^{d-1}\oplus\mathbb{R}$. In our case, however, we are interested in the partition function $\mathcal{Z}[\check{A}]$ defined over the entire spacetime $X^{d}$.
@@
There is indeed a way to interpret our vanishing result for the partition function in terms of the quantum Gauss law. It has become a standard understanding in the modern literature on generalised symmetries that global symmetries in a $d$-dimensional QFT are associated with edge modes of gauge symmetries in a $(d+1)$-dimensional Symmetry Topological Field Theory (SymTFT) Apruzzi:2021nmk ; Freed:2022qnc . From the perspective of the SymTFT, the partition function $\mathcal{Z}[\check{A}]$ of the boundary theory supported on $X^{d}$ is identified as a wavefunction $\Psi(\underline{\check{A}}|_{X})$ of the bulk theory, satisfying

### v3

@@ 2.2 Anomaly and the mapping cylinder
where $\mathcal{A}[\check{A},\check{\lambda}]$ is known as the ’t Hooft anomaly ’t Hooft (1980). For simplicity, we have assumed here that the partition function is a section of a line bundle $\mathcal{L}$ (i.e. a vector bundle with rank $1$) over the moduli space of gauge fields, $\mathfrak{A}\simeq\check{Z}^{p+2}(X^{d})$. Equivalently, $\mathcal{A}[\check{A},\check{\lambda}]$ can be realised in terms of an invertible field theory on a $(d+1)$-dimensional bulk $Y^{d+1}$ such that $\partial Y^{d+1}=X^{d}$. More generally, we can consider replacing the line bundle with a vector bundle of higher rank, but we will not study such cases in our work.
When $\mathcal{A}[\check{A},\check{\lambda}]\notin\mathbb{Z}$, one cannot uniquely define $\mathcal{Z}[\check{A}]$ for a given differential cohomology class $\check{A}\in\check{H}^{p+2}(X^{d})$. This means that the partition function is not well-defined over $\mathfrak{A}/\mathfrak{G}\simeq\check{H}^{p+2}(X^{d})$, i.e. the space of gauge fields modulo gauge transformations, where $\mathfrak{G}\simeq\check{C}^{p+1}_{\text{flat}}(X^{d})$, the space of degree-$(p+1)$ differential cochains with vanishing curvature component. Hence, the anomaly can be interpreted as an obstruction to gauging the global symmetry, i.e. summing over all inequivalent classes of background gauge fields $\check{A}\in\check{H}^{p+2}(X^{d})$ in the partition function. (If we insist in doing such a sum, the result will vanish, but this is a different kind of vanishing to the one in this paper, which does not involve gauging of the symmetry and happens only for specific background field configurations.) We stress that this notion of an anomaly applies not only to a $p$-form $\mathrm{U}(1)$ symmetry, but also to when the symmetry is any group, e.g. non-Abelian or discrete Hopkins and Singer (2005); Hsieh et al. (2022), and even when it is a higher group, as long as we define the suitable differential cocycles $\check{A}$ accordingly. In the latter case, $\mathcal{A}[\check{A},\check{\lambda}]$ is usually known as a mixed anomaly between the different levels of the associated Postnikov tower.
Geometrically, the relation (26) can be reinterpreted as a parallel transport of $\mathcal{Z}[\check{A}]$ over the mapping cylinder $X^{d}\times I$, where $I$ is the unit interval, which we parametrise with a coordinate $t\in[0,1]$. More explicitly, we formally extend the gauge field to $X^{d}\times I$ such that it interpolates between two different gauge fields $\check{A},\check{A}^{\prime}\in\check{Z}^{p+2}(X^{d})$ on the two ends (we will soon impose that $A$ and $A^{\prime}$ are related by a gauge transformation), i.e.
@@
where $\check{\mathbb{A}}\in\check{C}^{p+2}(X^{d}\times I)$ is a “differential cochain” represented by the triplet, [^6] The combination $(1-t)\mathsf{C}_{p+2}+t\mathsf{C}_{p+2}^{\prime}$ is generally not an integral cochain in $C^{p+2}(X^{d}\times I;\mathbb{Z})$ for arbitrary $t\in[0,1]$, so $\check{\mathbb{A}}$, and hence the bulk gauge field $\underline{\check{A}}$, are not strictly differential cochains as we defined in (11). However, we will ultimately be interested only in the case where $\mathsf{C}_{p+2}=\mathsf{C}_{p+2}^{\prime}$, so the $t$ dependence in the characteristic class of $\underline{\check{A}}$ will drop off.
@@
for some $\check{a}=(\mathsf{c}_{p+1},a_{p},f_{p+1})\in\check{C}^{p+1}(X^{d})$. [^7] To be precise, we should send both $\check{a}$ and $\check{t}$ to differential cochains in $\check{C}^{\ast}(X^{d}\times I)$ via the inclusion maps $\iota_{X}:X^{d}\hookrightarrow X^{d}\times I$ and $\iota_{I}:I\hookrightarrow X^{d}\times I$, in order to make sense of their product. When one sets $\check{a}=0$, the differential cochain $\underline{\check{A}}=\check{\mathbb{A}}\in\check{C}^{p+2}(X^{d})$ becomes “pure-shift” Simons and Sullivan (2008).
@@
In any given QFT, we shall define the anomaly $\mathcal{A}[\check{A},\check{\lambda},\check{a}]\in\mathbb{R}$ as in (26), but with an auxiliary gauge field $\check{a}$ included, as the evaluation of some rational polynomial $\mathcal{P}\in\mathbb{Q}[\check{A},\check{\lambda},\check{a},\check{t}]$ over the mapping cylinder $X^{d}\times I$. We will discuss some explicit examples of the polynomial $\mathcal{P}$ in the sections that follow. Recall that the bulk gauge field $\underline{\check{A}}\in\check{Z}^{p+2}(X^{d}\times I)$ is defined solely by the constraints,
@@ 2.3 Vanishing theorem for the partition function
for some choice of $\check{\lambda}\in\check{Z}_{\text{flat}}^{p+1}(X^{d})$ now regarded instead as an auxiliary gauge field. Since $\check{a}\in\check{Z}_{\text{flat}}^{p+1}(X^{d})$ is closed under the differential $d$ by construction, we have $\check{A}-d\check{a}=\check{A}$. In other words, the space of flat differential cocycles $\check{Z}_{\text{flat}}^{p+1}(X^{d})$ can be interpreted as the space of rigid gauge transformations that does not change the gauge field $\check{A}\in\check{Z}^{p+1}(X^{d})$. [^9] In the terminology of Diaconescu et al. (2003b); Moore (2005); Belov and Moore (2004, 2006); Freed et al. (2007), these were referred to as global gauge transformations. More explicitly, under such a transformation, we find
@@
As far as $\mathcal{A}[\check{A},0,\check{a}]$ is concerned, it will be convenient for us to carry out computations on the mapping torus $X^{d}\times S^{1}$ instead by gluing the two ends of the interval $I$, on which the bulk gauge field $\iota_{X\times\{0,1\}}^{\ast}\,\underline{\check{A}}=\check{A}\in\check{Z}^{p+2}(X^{d})$ becomes identical when pulled back to $X^{d}$. In fact, $\mathcal{A}[\check{A},0,\check{a}]$ can now be interpreted as a holonomy of the polynomial $\mathcal{P}$ over $X^{d}\times S^{1}$. The practical advantage of doing so is that the mapping torus has no boundary, so if needed $\mathcal{P}$ can be further promoted to an invariant polynomial having support over a bounding manifold $Z^{d+2}$ such that $\partial Z^{d+2}=X^{d}\times S^{1}$. [^10] In general, such an extension exists if and only if the pair $(X^{d}\times S^{1},\mathsf{C}_{p+2})$ is trivial as an element of the bordism group $\Omega_{d+1}(K(\mathbb{Z},p+2))$, where $K(\mathbb{Z},p+2)$ is an Eilenberg-MacLane space whose only non-trivial homotopy group is $\pi_{p+2}=\mathbb{Z}$. The bordism group should be replaced by its suitable variants if we wish the extension to preserve some extra tangential structures. For instance, it is common to require the extension to preserve the Spin structure, in which case $S^{1}$ is null-bordant if and only if it has a Neveu-Schwarz, i.e. bounding, Spin structure. For example, we can take $Z^{d+2}=X^{d}\times D^{2}$, where $D^{2}$ is the solid 2-disk with $\partial D^{2}=S^{1}$. More precisely, we would like to construct a differential cocycle $\check{\mathcal{P}}\in\check{H}^{d+2}(Z^{d+2})$ whose holonomy is given by $\chi(Z^{d+2})=\exp(2\pi i\int_{X^{d}\times S^{1}}\mathcal{P})=\mathcal{A}[\check{A},0,\check{a}]$. See Figure 1 for an illustration.
@@
For non-flat $\check{a}\in\check{Z}^{p+1}(X^{d})$, there is no obvious interpretation in terms of gauge transformations of the background gauge field $\check{A}\in\check{Z}^{p+2}(X^{d})$ in the $d$-dimensional QFT on $X^{d}$. One may naïvely try to use the correspondence (41), and regard $\check{a}$ as a local gauge transformation of $\check{A}$. However, this is not consistent with the fact that, by definition, a gauge equivalence is one that leaves the holonomy $\chi$ of a differential cocycle (in this case the anomaly polynomial $\mathcal{P}$) invariant, which corresponds to differential cochains, of one degree lower, that have vanishing curvature. This contradicts the assumption that $\check{a}$ is not flat. Henceforth, we restrict the auxiliary gauge field $\check{a}\in\check{Z}^{p+1}(X^{d})$ to be flat. [^11] See Diaconescu et al. (2003b); Belov and Moore (2004) for a related discussion of the more general case when $\check{a}$ is not flat. As a side remark, the requirement of $\check{a}$ being flat formally amounts to imposing a weight filtration of a differential function, here taken to be the bulk gauge field $\underline{\check{A}}$, as defined by Hopkins and Singer (2005) (see also Diaconescu et al. (2003b) for a concise review). However, we will not make use of this language for the purpose of our work.
@@
In contrast, the latter arises when one simply sits at some fixed basepoint $\check{A}\in\mathfrak{A}$, so a non-trivial phase $e^{2\pi i\mathcal{A}[\check{A},0,\check{a}]}$ signifies that the section $\mathcal{Z}[\check{A}]$ itself has to vanish at this point, as depicted in Figure 2. This happens because $\mathfrak{G}$ does not act freely, and its fixed points correspond precisely to the space of rigid gauge transformations, $\check{Z}_{\text{flat}}^{p+1}(X^{d})$. [^12] It is perhaps better to rephrase this picture in the language of moduli stacks. Loosely speaking, we should replace the moduli space with a lasagna-like object where each layer is a copy of $\mathfrak{A}$, then a rigid gauge transformation should correspond to a non-trivial path across layers while fixing the basepoint $\check{A}$ on each copy. This implies that the partition function $\mathcal{Z}[\check{A}]$, despite the notation, depends not only on the background gauge field, but also secretly on its automorphisms. A more detailed investigation on this perspective is left to future work. We also thank Victor Carmona for a useful discussion on this issue. For this reason, we will hereafter call $\mathcal{A}[\check{A},0,\check{a}]$ the basepoint anomaly. See Seiberg et al. (2018); Meynet et al. (2025) for a similar discussion in the context of anomalies of duality groups, in which case the “gauge field” is the coupling constant $\tau$, and $\mathfrak{A}/\mathfrak{G}$ becomes $\mathbb{H}/\text{SL}(2,\mathbb{Z})$, where $\mathbb{H}$ denotes the upper-half plane of $\mathbb{C}$.
Note that when $\mathcal{Z}[\check{A}]\neq 0$, the quantity $e^{2\pi i\mathcal{A}[\check{A},\check{\lambda},\check{a}]}$ satisfies a factorization property as follows. Compatibility with composition requires
@@
On the contrary, if $\mathcal{Z}[\check{A}]=0$, then it follows from (49) that $\mathcal{Z}[\check{A}-d\check{\lambda}]$ also vanishes. Therefore, the vanishing of the partition function is a gauge-invariant statement.
2.4 Relation to quantum Gauss law and SymTFT
In the context where the symmetry in question is not a global symmetry but a gauge symmetry of the QFT, our analysis reproduces a previous result in the literature concerning the gauge invariance of physical wavefunctions Diaconescu et al. (2003b); Moore (2005); Belov and Moore (2004, 2006); Freed et al. (2007) (see also Hosseini et al. (2025) for a recent study of anomaly cancellation in Type I string theory). Here, $\mathcal{L}$ is assumed to lift to a trivial line bundle over $\mathfrak{A}/\mathfrak{G}$, so the partition function is gauge-invariant, i.e. $\mathcal{Z}[\check{A}-d\check{\lambda}]=\mathcal{Z}[\check{A}]$, and generally non-vanishing. Suppose we first neglect the auxiliary gauge field $\check{a}\in\check{Z}_{\text{flat}}^{p+1}(X^{d})$, then we demand
@@ 2.4 Relation to quantum Gauss law and SymTFT
which is known as the classical Gauss law. [^14] At a semi-classical level, Gauss law constraints can be used to construct topological operators generating the respective global symmetries. See e.g. Apruzzi et al. (2023a); Choi et al. (2023b); Bah et al. (2024) for some recent applications to non-invertible symmetries. On the contrary, one should also impose that the wavefunction is invariant under the action of non-trivial automorphisms $\check{a}$ of $\check{A}$. This further refines the constraint above, i.e.
@@
More precisely, despite giving similar conclusions, there is a subtle conceptual difference between Diaconescu et al. (2003b); Moore (2005); Belov and Moore (2004, 2006); Freed et al. (2007) and our approach. In the former, one is interested in wavefunctions $\Psi(\check{A}|_{W})$ obtained after performing a Hamiltonian quantisation on the theory. This effectively treats $X^{d}$ as a manifold with boundary $\partial X^{d}=W^{d-1}$, such that the tangent bundle near the “time-slice” $W^{d-1}$ decomposes as $TX^{d}|_{W}\cong TW^{d-1}\oplus\mathbb{R}$. In our case, however, we are interested in the partition function $\mathcal{Z}[\check{A}]$ defined over the entire spacetime $X^{d}$.
@@
There is indeed a way to interpret our vanishing result for the partition function in terms of the quantum Gauss law. It has become a standard understanding in the modern literature on generalised symmetries that global symmetries in a $d$-dimensional QFT are associated with edge modes of gauge symmetries in a $(d+1)$-dimensional Symmetry Topological Field Theory (SymTFT) Apruzzi et al. (2023b); Freed et al. (2022). The anomaly theory is, loosely speaking, the SymTFT without the dynamics. For this particular discussion, it is more convenient to treat the latter (rather than the former) as a gauge theory in its own right, so that it makes more sense to talk about Gauss law constraints here. From the perspective of the SymTFT, the partition function $\mathcal{Z}[\check{A}]$ of the boundary theory supported on $X^{d}$ is identified as a wavefunction $\Psi(\underline{\check{A}}|_{X})$ of the bulk theory, satisfying